Hi,
I'm having trouble in a section where we are to graph a parabola with horizontal and vertical axis. For example, the problem x=y^2-10y+30. I got the vertex (5,5) and since a=1 I know it opens to the right. I know for any constant d, the points on the parabola with y-coordinates 5-d and 5+d have the same x-coordinate. So I plug in 1 and have the y coordinates as ( ,4) and ( ,6). My problem is, I can't find the constant x-coordinate. I don't know how to get the x-coordinate. Can someone please help me figure out how to find the x-coordinate in this problem. I've tried everything and can't figure it out. Thanks so much!

Answer by stanbon(48570)

For example, the problem x=y^2-10y+30. I got the vertex (5,5) and since a=1 I know it opens to the right. I know for any constant d, the points on the parabola with y-coordinates 5-d and 5+d have the same x-coordinate. So I plug in 1 and have the y coordinates as ( ,4) and ( ,6). My problem is, I can't find the constant x-coordinate. I don't know how to get the x-coordinate. Can someone please help me figure out how to find the x-coordinate in this problem.
-------------
You have the y-coordinate of two points.
If y=4, x=4^2-10*4+30; x=46-40=6
If y=6, x=6^2-10*6+30; x=66-60=6
-----------
For both of yur points the x-value is 6.
==============
Cheers,
Stan H.